Micro Spinoffs Inc. issued 10-year debt a year ago at par value with a coupon rate of 8%, paid annually. Today, the debt is selling at $1,280. If the firm’s tax bracket is 21%, what is its percentage cost of debt? Assume a face value of $1,000. (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Answer #1

Annual coupon = 0.08 * 1000 = 80

YTM = 4.1998%

Keys to use in a financial calculator:

FV 1000

PV -1280

PMT 80

N 9

CPT I/Y

Cost of debt = YTM (1 - tax)

Cost of debt = 0.041998 (1 - 0.21)

Cost of debt = 0.0332 or 3.32%